# Operator World Models for Reinforcement Learning

### v1

@@ Abstract
Policy Mirror Descent (PMD) is a powerful and theoretically sound methodology for sequential decision-making. However, it is not directly applicable to Reinforcement Learning (RL) due to the inaccessibility of explicit action-value functions. We address this challenge by introducing a novel approach based on learning a world model of the environment using conditional mean embeddings. We then leverage the operatorial formulation of RL to express the action-value function in terms of this quantity in closed form via matrix operations. Combining these estimators with PMD leads to POWR, a new RL algorithm for which we prove convergence rates to the global optimum. Preliminary experiments in finite and infinite state settings support the effectiveness of our method.
@@ 1 Introduction
In this work we address these issues by proposing a novel approach to estimating the action-value function. Unlike previous methods that directly approximate the action-value function from samples, we first learn the transition operator and reward function associated with the Markov decision process (MDP). To model the transition operator, we adopt the Conditional Mean Embedding (CME) framework [14, 15]. We then leverage an operatorial characterization of the action-value function to express it in terms of these estimated quantities. This strategy draws a peculiar connection with world model methods [16], and can be interpreted as world model-learning via CMEs. However, traditional world model methods emphasize learning an implicit model of the environment in the form of a simulator. This requires extensive sampling for application to PMD and incurs into to two sources of error in estimating the action-value function: model and sampling error. In contrast, CMEs can be used to estimate expectations without sampling and incur only in model error, for which learning bounds are available [17, 18]. One of our key results shows that by modeling the transition operator as a CME between suitable Sobolev spaces, we can compute estimates of the action-value function of any sufficiently smooth policy in closed form via efficient matrix operations.
Combining our estimates of the action-value function with the PMD framework we obtain a novel RL algorithm that we dub Policy mirror descent with Operator World-models for Reinforcement learning (POWR). A byproduct of adopting CMEs to model the transition operator is that we can naturally extend PMD to infinite state space settings. We leverage recent advancements in characterizing the sample complexity of CME estimators to prove convergence rates for the proposed algorithm to the global maximum of the RL Problem. Our approach is similar in spirit to [19], which proposed a policy optimization strategy based on CMEs. We extend these ideas to PMD strategies and refine previous results on convergence rates as a byproduct of our analysis. Learning the transition operator with a least-squares based estimator was also recently considered in [20], which proposed an optimistic strategy to prove near-optimal regret bounds in linear mixture MDP settings [21]. In contrast, in this work we cast our problem within an linear MDP setting with possibly infinite latent dimension. We validate empirically our approach in both finite and infinite state settings, reporting promising evidence in support of our theoretical analysis.
@@
The main contributions of this paper are: $i)$ a CME-based world model estimator, which enables us to generate estimators for the action-value function of a policy in closed form via matrix operations. $ii)$ An (inexact) PMD algorithm combining the learned CMEs world models with mirror descent update steps to generate improved policies. $iii)$ Showing that the algorithm is well-defined when learning the world model as an operator between a suitable family of Sobolev spaces. $iv)$ Showing convergence rates of the proposed approach to the global maximum of the RL problem, under regularity assumptions on the MDP. $v)$ Empirically testing the proposed approach in practice, comparing it with well-established baselines.
@@ 2 Problem Formulation and Policy Mirror Descent
We consider a Markov Decision Process (MDP) over a state space $\mathcal{X}$ and action space $\mathcal{A}$, with transition kernel $\tau$. We assume $\mathcal{X}$ and $\mathcal{A}$ to be Polish, $\tau:\Omega\to{\mathcal{M}}_{1}^{+}(\mathcal{X})$ to be a Borel measurable function from the joint space $\Omega=\mathcal{X}\times\mathcal{A}$ to the space ${\mathcal{M}}_{1}^{+}(\mathcal{X})$ of Borel probability measures on ${\mathcal{X}}$. We define a policy to be a Borel measurable function $\pi:{\mathcal{X}}\to{\mathcal{M}}_{1}^{+}(\mathcal{A})$. When $\mathcal{A}$ (respectively $\mathcal{X}$) is a finite set, the space ${\mathcal{M}}_{1}^{+}(\mathcal{A})=\Delta(\mathcal{A})\subseteq\mathbb{R}^{|A|}$ (respectively ${\mathcal{M}}_{1}^{+}(\mathcal{X})=\Delta(\mathcal{X})\subseteq\mathbb{R}^{|X|}$) corresponds to the probability simplex. Given a discount factor $\gamma>0$, a (starting) state distribution $\nu\in{\mathcal{M}}_{1}^{+}(\mathcal{X})$ and a Borel measurable bounded and non-negative reward [^1] All the discussion in this work can be extended to the case where also the rewards are random and $\tau:\Omega\to{\mathcal{M}}_{1}^{+}({\mathcal{X}}\times\mathbb{R})$ takes values in the space of joint distributions over states and rewards $(X_{t+1},R_{t})$ function $r:\Omega\to\mathbb{R}$ we denote by
$J(\pi)=E\left[\sum_{t=0}^{\infty}\gamma^{t}r(X_{t},A_{t})\right]$ (1)
@@
In sequential decision settings, the goal is to find the optimal policy $\pi_{*}$ maximizing 1 over the space of all measurable policies. In reinforcement learning, one typically assumes that knowledge of the transition $\tau$, the reward $r$, and (possibly) the starting distribution $\nu$ is not available. It is only possible to gather information about these quantities by interacting with the MDP to sample trajectories of state-action pairs $(x_{t},a_{t})$ and corresponding rewards $r(x_{t},a_{t})$.
@@
In so-called tabular settings – in which both $\mathcal{X}$ and $\mathcal{A}$ are finite sets – the policy optimization problem amounts to maximizing 1 over the space $\Pi=\Delta(\mathcal{A})\otimes\mathbb{R}^{|\mathcal{X}|}$ of column substochastic matrices, namely matrices $M\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{X}|}$ with non-negative entries and whose columns sum up to one, namely $M^{*}\mathbf{1}_{\mathcal{A}}=\mathbf{1}_{\mathcal{X}}$, with $\mathbf{1}$ denoting the vector with all entries equal to one on the appropriate space. Borrowing from the convex optimization literature – where mirror descent algorithms offer a powerful approach to minimize a convex functional over a convex constraint set [22, 23] – recent work proposed to adopt mirror descent also for policy optimization, a strategy known as policy mirror descent (PMD) [10]. Even though the objective in 1 is not convex (or concave, since we are maximizing it [11]), it turns out that mirror ascent can nevertheless enjoy global convergence to the maximum, with sublinear [11] or even linear rates [12] (at the cost of dimension-dependent constants).
Starting from an initial policy $\pi_{0}$, PMD generates a sequence $(\pi_{t})_{t\in\mathbb{N}}$ according to the the update step
$\pi_{t+1}(\cdot\,|\,x)=\operatorname*{argmin}_{p\in\Delta(\mathcal{A})}\quad-\eta\left\langle{q_{\pi_{t}}(\cdot,x)},{p}\right\rangle+D(p,\pi_{t}(\cdot|x)),$ (2)
for any $x\in\mathcal{X}$, with $\eta>0$ a step size, $D$ a suitable Bregman divergence [23] and $q_{\pi}:\Omega\to\mathbb{R}$ the so-called action-value function of a policy $\pi$, namely
$q_{\pi}(x,a)=E\left[\sum_{t=0}^{\infty}\gamma^{t}r(X_{t},A_{t})\middle|X_{0}=x,A_{0}=a\right],$ (3)
the discounted return obtained by taking action $a\in\mathcal{A}$ in state $x\in\mathcal{X}$ and then following the policy $\pi$. The solution to 2 crucially depends on the choice of $D$. For example, in [10] the authors observed that if $D$ is the Kullback-Leibler divergence, PMD corresponds to the Natural Policy Gradient originally proposed in [24] while [12] showed that if $D$ is the squared euclidean distance, PMD recovers the Projected Policy Gradient method from [11].
@@
The work in [12] provides an answer to this question, extending the analysis of PMD to the case where estimates $\hat{q}_{t}$ are used in in place of the true action-value function in 2. We recall here an informal version of the result for the case of sublinear convergence rates for PMD. We postpone a more rigorous statement of the theorem and its assumptions to Sec. 5, where we extend it to infinite state spaces $\mathcal{X}$.
@@
In the tabular setting, let $(\pi_{t})_{t\in\mathbb{N}}$ be a sequence of policies obtained by applying the PMD update in 2 with functions $\hat{q}_{t}:\Omega\to\mathbb{R}$ in place of $q_{\pi_{t}}$ and $D$ a suitable Bregman divergence. For any $T\in\mathbb{N}$ and $\varepsilon>0$, if $\left\|{\hat{q}_{t}-q_{\pi_{t}}}\right\|_{\infty}\leq\varepsilon$ for all $t=1,\dots,T$, then
@@
Thm. 1 implies that inexact PMD retains the convergence rates of its exact counterpart, provided that the approximation error for each action-value function is of order $1/T$ in uniform norm $\left\|{\cdot}\right\|_{\infty}$. While this result supports estimating the action-value function in RL, implementing this strategy in practice poses two main challenges, even in tabular settings. First, approximating the expectation in 3 in $\left\|{\cdot}\right\|_{\infty}$ norm via sampling requires “starting” the MDP from each state $x\in\mathcal{X}$, multiple times. This is often not possible in RL, where we do not have control over the starting distribution $\nu$. Second, repeating this sampling process to learn a $\hat{q}_{t}$ for each policy $\pi_{t}$ can become extremely expensive in terms of both the number of computations and interactions with the environment.
In this work we propose a new strategy to tackle the problems above. Rather than sampling the MDP to directly estimate each $\hat{q}_{t}$ we learn estimators $\hat{r}$ and $\hat{\tau}$ for the reward and transition distribution respectively. Then, for any policy $\pi$, we leverage the relation between these quantities in 3 to generate an estimator $\hat{q}$ for $q_{\pi}$. This tackles the above challenges since: 1) it enables us to control the approximation error on any action-value function in terms the approximation error of $\hat{r}$ and $\hat{\tau}$; 2) it does not require sampling from the MDP to learn a new $\hat{q}_{t}$ for each $\pi_{t}$ generated by PMD.
3 Proposed Estimator and Algorithm
In this section, we consider the operator-based formulation of the problem introduced in Sec. 2 (see also [11]). This will be instrumental to extend the PMD theory to arbitrary state spaces $\mathcal{X}$, to quantify the relation between the approximation error of the action-value in terms of the approximation error of the reward and transition distribution, and to motivate conditional mean embeddings as the tool to learn these latter quantities.
@@ 3 Proposed Estimator and Algorithm
We start by defining the transition operator ${T}$ associated to the MDP transition distribution $\tau$. Let $B_{b}(\mathcal{X})$ denote the space of bounded Borel measurable functions on a space $\mathcal{X}$. Formally, ${T}:B_{b}(\mathcal{X})\to B_{b}(\Omega)$ is the linear operator such that, for any $f\in B_{b}(\mathcal{X})$
$({T}f)(x,a)=\int_{\mathcal{X}}f(x^{\prime})\leavevmode\nobreak\ \tau(dx^{\prime}|x,a)=E\left[f(X^{\prime})\mid x,a\right]\qquad\text{for all }(x,a)\in\Omega,$ (5)
where $X^{\prime}$ is sampled according to $\tau(\cdot|x,a)$. Note that ${T}$ is the Markov operator [25, Ch. 19] encoding the dynamics of the MDP and its conjugate ${T}^{*}:\mathcal{M}(\Omega)\to\mathcal{M}({\mathcal{X}})$ is the operator mapping measures $\mu\in\mathcal{M}(\Omega)$ to their transition via $\tau$ to $(T^{*}\mu)(\mathcal{B})=\int_{\mathcal{B}\times\Omega}\tau(dx^{\prime}|x,a)\mu(dx,da)$ for any measurable $\mathcal{B}\subseteq\mathcal{X}$. For any policy $\pi$ we define the operator ${P}_{\pi}:B_{b}(\Omega)\to B_{b}(\mathcal{X})$ such that for all $g\in B_{b}(\Omega)$
$({P}_{\pi}g)(x)=\int_{\mathcal{A}}g(x,a)\leavevmode\nobreak\ \pi(da|x)=E\left[g(X,A)\mid X=x\right]\quad\text{for all }x\in\mathcal{X},$ (6)
where the expectation is taken over the action $A$ sampled according to $\pi(\cdot|x)$. Also ${P}_{\pi}$ is a Markov operator and its conjugate ${P}_{\pi}^{*}:\mathcal{M}(\mathcal{X})\to\mathcal{M}(\Omega)$ is the operator mapping any $\nu\in\mathcal{M}({\mathcal{X}})$ to its “multiplication” by $\pi$, namely $({P}_{\pi}^{*}\nu)(\mathcal{C})=\int_{\mathcal{C}}\pi(da|x)\nu(dx)$ for any measurable $\mathcal{C}\subseteq\Omega$.
@@
With these two operators in place, we can characterize the expected reward after a single interaction between a policy $\pi$ and the MDP as $({T}{P}_{\pi}r)(x,a)=E[r(X^{\prime},A^{\prime})|X_{0}=x,A_{0}=a]$. This observation can be applied recursively, yielding the operatorial characterization of the action-value function from 3
$q_{\pi}(x,a)=\sum_{t=0}^{\infty}\gamma^{t}E[r(X_{t},A_{t})|X_{0}=x,A_{0}=a]=\sum_{t=0}^{\infty}(\gamma{T}{P}_{\pi})^{t}r=({Id}-\gamma{T}{P}_{\pi})^{-1}r,$ (7)
where the last equality follows from ${T}$ and ${P}_{\pi}$ being Markov operators [25, Ch. 19] ($\left\|{{T}}\right\|=\left\|{{P}_{\pi}}\right\|=1$), making the Neumann series convergent. Analogously, we can reformulate the RL objective introduced in 1 as the pairing
$J(\pi)=\left\langle{{P}_{\pi}({Id}-\gamma{T}{P}_{\pi})^{-1}r},{\nu}\right\rangle=\left\langle{{P}_{\pi}q_{\pi}},{\nu}\right\rangle,$ (8)
for $\nu\in{\mathcal{M}}_{1}^{+}(\mathcal{X})$ a starting distribution. In both 7 and 8 the operatorial formulation encodes the cumulative reward collected through the (possibly infinitely many) interactions of the policy with the MDP in closed form, as the inversion $({Id}-\gamma{T}{P}_{\pi})^{-1}r$. This characterization motivates us to learn ${T}$ and $r$ from data and then express any action-value function as the interaction of these two terms with the policy $\pi$ as in 7, rather than learning each $q_{\pi}$ independently for any $\pi$.
@@
Conditional Mean Embeddings (CME) offer an effective tool to model and learn conditional expectation operators from data [15]. They cast the problem of learning ${T}$ by studying the restriction of its action on a suitable family of functions. Let $\varphi:\mathcal{X}\to{\mathcal{F}}$ and $\psi:\Omega\to{\mathcal{G}}$ two feature maps with values into the Hilbert spaces ${\mathcal{F}}$ and ${\mathcal{G}}$. With some abuse of notation (which is justified by them being Hilbert spaces), we interpret ${\mathcal{F}}$ and ${\mathcal{G}}$ as subspaces of functions in $B_{b}(\mathcal{X})$ and $B_{b}(\Omega)$ of the form $f(x)=\left\langle{f},{\varphi(x)}\right\rangle$ and $g(x,a)=\left\langle{g},{\psi(x,a)}\right\rangle$ for any $f\in{\mathcal{F}}$ and $g\in{\mathcal{G}}$ and any $(x,a)\in\Omega$. We say that the linear MDP assumption holds with respect to $(\varphi,\psi)$ if
@@
The restriction of  ${T}$ to ${\mathcal{F}}$ is a Hilbert-Schmidt operator ${T}|_{\mathcal{F}}\in{HS}({\mathcal{F}},{\mathcal{G}})$.
In CME settings, the assumption above is known as requiring the CME of $\tau$ to be well-specified. The following result clarifies this aspect, as well as establishing the relation of Asm. 1 with the standard definition of linear MDP.

### v2

@@ Abstract
Policy Mirror Descent (PMD) is a powerful and theoretically sound methodology for sequential decision-making. However, it is not directly applicable to Reinforcement Learning (RL) due to the inaccessibility of explicit action-value functions. We address this challenge by introducing a novel approach based on learning a world model of the environment using conditional mean embeddings. Leveraging tools from operator theory we derive a closed-form expression of the action-value function in terms of the world model via simple matrix operations. Combining these estimators with PMD leads to POWR, a new RL algorithm for which we prove convergence rates to the global optimum. Preliminary experiments in finite and infinite state settings support the effectiveness of our method [^3] Code available at: github.com/CSML-IIT-UCL/powr.
@@ 1 Introduction
In this work, we address these issues by proposing a novel approach to estimating the action-value function. Unlike previous methods that directly approximate the action-value function from samples, we first learn the transition operator and reward function associated with the Markov decision process (MDP). To model the transition operator, we adopt the Conditional Mean Embedding (CME) framework [14, 15]. We then leverage an operatorial characterization of the action-value function to express it in terms of these estimated quantities. This strategy draws a peculiar connection with world model methods and can be interpreted as world model learning via CMEs. The notion of world models for RL has been popularized by Ha and Schmidhuber in [16] distilling ideas from the early nineties [17, 18]. Traditional world model methods such as [16, 19] emphasize learning an implicit model of the environment in the form of a simulator. The simulator can be sampled directly in the latent representation space, which is usually of moderate dimension, resulting in a compressed and high-throughput model of the environment. This approach, however, requires extensive sampling for application to PMD and incurs into two sources of error in estimating the action-value function: model and sampling error. In contrast, CMEs can be used to estimate expectations without sampling and incur only in model error, for which learning bounds are available [20, 21]. One of our key results shows that by modeling the transition operator as a CME between suitable Sobolev spaces, we can compute estimates of the action-value function of any sufficiently smooth policy in closed form via efficient matrix operations.
Combining our estimates of the action-value function with the PMD framework we obtain a novel RL algorithm that we dub Policy mirror descent with Operator World-models for Reinforcement learning (POWR). A byproduct of adopting CMEs to model the transition operator is that we can naturally extend PMD to infinite state space settings. We leverage recent advancements in characterizing the sample complexity of CME estimators to prove convergence rates for the proposed algorithm to the global maximum of the RL Problem. Our approach is similar in spirit to [22], which proposed a value iteration strategy based on CMEs. We extend these ideas to PMD strategies and refine previous results on convergence rates. Learning the transition operator with a least-squares based estimator was also recently considered in [23], which proposed an optimistic strategy to prove near-optimal regret bounds in linear mixture MDP settings [24]. In contrast, in this work, we cast our problem within a linear MDP setting with possibly infinite latent dimension. We validate our approach on simple environments from the Gym library [25] both in finite and infinite state settings, reporting promising evidence in support of our theoretical analysis.
@@
The main contributions of this paper are: $i)$ a CME-based world model framework, which enables us to generate estimators for the action-value function of a policy in closed form via matrix operations. $ii)$ An (inexact) PMD algorithm combining the learned CMEs world models with mirror descent update steps to generate improved policies. $iii)$ Showing that the algorithm is well-defined when learning the world model as an operator between a suitable family of Sobolev spaces. $iv)$ Showing convergence rates of the proposed approach to the global maximum of the RL problem, under regularity assumptions on the MDP. $v)$ Empirically testing the proposed approach in practice, comparing it with well-established baselines.
@@ 2 Problem Formulation and Policy Mirror Descent
We consider a Markov Decision Process (MDP) over a state space $\mathcal{X}$ and action space $\mathcal{A}$, with transition kernel $\tau$. We assume $\mathcal{X}$ and $\mathcal{A}$ to be Polish, $\tau:\Omega\to\mathcal{P}(\mathcal{X})$ to be a Borel measurable function from the joint space $\Omega=\mathcal{X}\times\mathcal{A}$ to the space $\mathcal{P}(\mathcal{X})$ of Borel probability measures on ${\mathcal{X}}$. We define a policy to be a Borel measurable function $\pi:{\mathcal{X}}\to\mathcal{P}(\mathcal{A})$. When $\mathcal{A}$ (respectively $\mathcal{X}$) is a finite set, the space $\mathcal{P}(\mathcal{A})=\Delta(\mathcal{A})\subseteq{\mathbb{R}}^{|A|}$ (respectively $\mathcal{P}(\mathcal{X})=\Delta(\mathcal{X})\subseteq{\mathbb{R}}^{|X|}$) corresponds to the probability simplex. Given a discount factor $\gamma>0$, an initial state distribution $\nu\in\mathcal{P}(\mathcal{X})$ and a Borel measurable bounded and non-negative reward [^4] All the discussion in this work can be extended to the case where also the rewards are random and $\tau:\Omega\to\mathcal{P}({\mathcal{X}}\times{\mathbb{R}})$ takes values in the space of joint distributions over states and rewards $(X_{t+1},R_{t})$ function $r:\Omega\to{\mathbb{R}}$ we denote by
$J(\pi)=\mathbb{E}_{\nu,\pi,\tau}\left[\sum_{t=0}^{\infty}\gamma^{t}r(X_{t},A_{t})\right]$ (1)
@@
In sequential decision settings, the goal is to find the optimal policy $\pi_{*}$ maximizing 1 over the space of all measurable policies. In reinforcement learning, one typically assumes that knowledge of the transition $\tau$, the reward $r$, and (possibly) the starting distribution $\nu$ is not available. It is only possible to gather information about these quantities by interacting with the MDP to sample state-action pairs $(x_{t},a_{t})$ and corresponding rewards $r(x_{t},a_{t})$ and transitions $x_{t+1}$.
@@
In so-called tabular settings – in which both $\mathcal{X}$ and $\mathcal{A}$ are finite sets – the policy optimization problem amounts to maximizing 1 over the space $\Pi=\Delta(\mathcal{A})\otimes{\mathbb{R}}^{|\mathcal{X}|}$ of column substochastic matrices, namely matrices $M\in{\mathbb{R}}^{|\mathcal{A}|\times|\mathcal{X}|}$ with non-negative entries and whose columns sum up to one, namely $M^{*}\mathbf{1}_{\mathcal{A}}=\mathbf{1}_{\mathcal{X}}$, with $\mathbf{1}$ denoting the vector with all entries equal to one on the appropriate space. Borrowing from the convex optimization literature – where mirror descent algorithms offer a powerful approach to minimize a convex functional over a convex constraint set [26, 27] – recent work proposed to adopt mirror descent also for policy optimization, a strategy known as policy mirror descent (PMD) [10]. Even though the objective in 1 is not convex (or concave, since we are maximizing it), it turns out that mirror ascent can nevertheless enjoy global convergence to the maximum, with sublinear [11] or even linear rates [12], at the cost of dimension-dependent constants.
Starting from an initial policy $\pi_{0}$, PMD generates a sequence $(\pi_{t})_{t\in{\mathbb{N}}}$ according to the update step
$\pi_{t+1}(\cdot\,|\,x)=\operatornamewithlimits{argmin}_{p\in\Delta(\mathcal{A})}\quad-\eta\left\langle{q_{\pi_{t}}(\cdot,x)},{p}\right\rangle+D(p,\pi_{t}(\cdot|x)),$ (2)
for any $x\in\mathcal{X}$, with $\eta>0$ a step size, $D$ a suitable Bregman divergence [27] and $q_{\pi}:\Omega\to{\mathbb{R}}$ the so-called action-value function of a policy $\pi$, see also [12, Sec. 4]. The action-value function
$q_{\pi}(x,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(X_{t},A_{t})\middle|X_{0}=x,A_{0}=a\right]$ (3)
is the discounted return obtained by taking action $a\in\mathcal{A}$ in state $x\in\mathcal{X}$ and then following the policy $\pi$. The solution to 2 crucially depends on the choice of $D$. For example, in [10] the authors observed that if $D$ is the Kullback-Leibler divergence, PMD corresponds to the Natural Policy Gradient originally proposed in [28] while [12] showed that if $D$ is the squared euclidean distance, PMD recovers the Projected Policy Gradient method from [11].
@@
The work in [12] provides an answer to this question, extending the analysis of PMD to the case where estimates $\hat{q}_{\pi_{t}}$ are used in place of the true action-value function in 2. We recall here an informal version of the result for the case of sublinear convergence rates for PMD. We postpone a more rigorous statement of the theorem and its assumptions to Sec. 5, where we extend it to infinite state spaces $\mathcal{X}$.
@@
In the tabular setting, let $(\pi_{t})_{t\in{\mathbb{N}}}$ be a sequence of policies obtained by applying the PMD update in 2 with functions $\hat{q}_{\pi_{t}}:\Omega\to{\mathbb{R}}$ in place of $q_{\pi_{t}}$ and $D$ a suitable Bregman divergence. For any $T\in{\mathbb{N}}$ and $\varepsilon>0$, if $\left\|{\hat{q}_{\pi_{t}}-q_{\pi_{t}}}\right\|_{\infty}\leq\varepsilon$ for all $t=1,\dots,T$, then
@@
Thm. 1 implies that inexact PMD retains the convergence rates of its exact counterpart, provided that the approximation error for each action-value function is of order $1/T$ in uniform norm $\left\|{\cdot}\right\|_{\infty}$. While this result supports estimating the action-value function in RL, implementing this strategy in practice poses two main challenges, even in tabular settings. First, approximating the expectation in 3 in $\left\|{\cdot}\right\|_{\infty}$ norm via sampling requires “starting” the MDP from each state $x\in\mathcal{X}$, multiple times. This is often not possible in RL, where we do not have control over the starting distribution $\nu$. Second, repeating this sampling process to learn a $\hat{q}_{\pi_{t}}$ for each policy $\pi_{t}$ can become extremely expensive in terms of both the number of computations and interactions with the environment.
In this work, we propose a new strategy to tackle the problems above. Instead of re-sampling the MDP to estimate $\hat{q}_{\pi_{t}}$ at each iteration $t$, we learn estimators $\hat{r}$ and $\hat{\tau}$ for the reward and transition distribution, respectively. For any policy $\pi$, we then leverage the relation between these quantities in 3 to generate an estimator $\hat{q}$ for $q_{\pi}$. This approach tackles the above challenges since 1) it enables us to control the approximation error on any action-value function in terms of the approximation error of $\hat{r}$ and $\hat{\tau}$; 2) it does not require sampling the MDP to learn a new $\hat{q}_{\pi_{t}}$ for each $\pi_{t}$ generated by PMD.
3 Operator World Models
In this section, we present an operator-based formulation of the problem introduced in Sec. 2 (see also [11]). This will be instrumental in extending the PMD theory to arbitrary state spaces $\mathcal{X}$, to quantify the approximation error of the action-value function in terms of the approximation error of the reward and transition distribution, and to motivate conditional mean embeddings as the tool to learn these latter quantities.
@@ 3 Operator World Models
We start by defining the transfer operator $\mathsf{T}$ associated with the MDP transition distribution $\tau$. Let $B_{b}(\mathcal{X})$ denote the space of bounded Borel measurable functions on a space $\mathcal{X}$. Formally, $\mathsf{T}:B_{b}(\mathcal{X})\to B_{b}(\Omega)$ is the linear operator such that, for any $f\in B_{b}(\mathcal{X})$
$(\mathsf{T}f)(x,a)=\int_{\mathcal{X}}f(x^{\prime})\leavevmode\nobreak\ \tau(dx^{\prime}|x,a)=\mathbb{E}\left[f(X^{\prime})\mid x,a\right]\qquad\text{for all }(x,a)\in\Omega,$ (5)
where $X^{\prime}$ is sampled according to $\tau(\cdot|x,a)$. Note that $\mathsf{T}$ is the Markov operator [29, Ch. 19] encoding the dynamics of the MDP and its conjugate $\mathsf{T}^{*}:\mathcal{M}(\Omega)\to\mathcal{M}({\mathcal{X}})$ is the operator mapping signed Borel measures $\mu\in\mathcal{M}(\Omega)$ to their transition via $\tau$ as $(\mathsf{T}^{*}\mu)(\mathcal{B})=\int_{\mathcal{B}\times\Omega}\tau(dx^{\prime}|x,a)\mu(dx,da)$ for any measurable $\mathcal{B}\subseteq\mathcal{X}$. For any policy $\pi$ we define the operator $\mathsf{P}_{\pi}:B_{b}(\Omega)\to B_{b}(\mathcal{X})$ such that for all $g\in B_{b}(\Omega)$
$(\mathsf{P}_{\pi}g)(x)=\int_{\mathcal{A}}g(x,a)\leavevmode\nobreak\ \pi(da|x)=\mathbb{E}\left[g(X,A)\mid X=x\right]\quad\text{for all }x\in\mathcal{X},$ (6)
where the expectation is taken over the action $A$ sampled according to $\pi(\cdot|x)$. Also $\mathsf{P}_{\pi}$ is a Markov operator and its conjugate $\mathsf{P}_{\pi}^{*}:\mathcal{M}(\mathcal{X})\to\mathcal{M}(\Omega)$ is the operator mapping any $\nu\in\mathcal{M}({\mathcal{X}})$ to its joint measure with $\pi$, namely $(\mathsf{P}_{\pi}^{*}\nu)(\mathcal{C})=\int_{\mathcal{C}}\pi(da|x)\nu(dx)$ for any measurable $\mathcal{C}\subseteq\Omega$.
@@
With these two operators in place, we can characterize the expected reward after a single interaction between a policy $\pi$ and the MDP as $(\mathsf{T}\mathsf{P}_{\pi}r)(x,a)=\mathbb{E}[r(X^{\prime},A^{\prime})|X_{0}=x,A_{0}=a]$. This observation can be applied recursively, yielding the operatorial characterization of the action-value function from 3
$q_{\pi}(x,a)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[r(X_{t},A_{t})|X_{0}=x,A_{0}=a]=\sum_{t=0}^{\infty}(\gamma\mathsf{T}\mathsf{P}_{\pi})^{t}r=(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{\pi})^{-1}r,$ (7)
where the last equality follows from $\mathsf{T}$ and $\mathsf{P}_{\pi}$ being Markov operators [29, Ch. 19] ($\left\|{\mathsf{T}}\right\|=\left\|{\mathsf{P}_{\pi}}\right\|=1$), making the Neumann series convergent. Analogously, we can reformulate the RL objective introduced in 1 as the pairing
$J(\pi)=\left\langle{\mathsf{P}_{\pi}(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{\pi})^{-1}r},{\nu}\right\rangle=\left\langle{\mathsf{P}_{\pi}q_{\pi}},{\nu}\right\rangle,$ (8)
for $\nu\in\mathcal{P}(\mathcal{X})$ a starting distribution. In both 7 and 8 the operatorial formulation encodes the cumulative reward collected through the (possibly infinitely many) interactions of the policy with the MDP in closed form, as the inversion $(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{\pi})^{-1}r$. This characterization motivates us to learn $\mathsf{T}$ and $r$ from data and then express any action-value function as the interaction of these two terms with the policy $\pi$ as in 7, rather than learning each $q_{\pi}$ independently for any $\pi$.
@@
Conditional Mean Embeddings (CME) offer an effective tool to model and learn conditional expectation operators from data [15]. They cast the problem of learning $\mathsf{T}$ by studying the restriction of its action on a suitable family of functions. Let $\varphi:\mathcal{X}\to{\mathcal{F}}$ and $\psi:\Omega\to{\mathcal{G}}$ two feature maps with values into the Hilbert spaces ${\mathcal{F}}$ and ${\mathcal{G}}$. With some abuse of notation (which is justified by them being Hilbert spaces), we interpret ${\mathcal{F}}$ and ${\mathcal{G}}$ as subspaces of functions in $B_{b}(\mathcal{X})$ and $B_{b}(\Omega)$ of the form $f(x)=\left\langle{f},{\varphi(x)}\right\rangle$ and $g(x,a)=\left\langle{g},{\psi(x,a)}\right\rangle$ for any $f\in{\mathcal{F}}$ and $g\in{\mathcal{G}}$ and any $(x,a)\in\Omega$. We say that the linear MDP assumption holds with respect to $(\varphi,\psi)$ if
@@
The restriction of  $\mathsf{T}$ to ${\mathcal{F}}$ is a Hilbert-Schmidt operator $\mathsf{T}|_{\mathcal{F}}\in\mathsf{HS}({\mathcal{F}},{\mathcal{G}})$.
In CME settings, the assumption above is known as requiring the CME of $\tau$ to be well-specified. The following result, proved in Sec. A.2, clarifies this aspect and establishes the relation of Asm. 1 with the standard definition of linear MDP.